18. Suppose R^2 has weighted inner product given as <u, v> =(3u base 1 v base 1 + 2u base 2 V base 2 for u = (u base 1, u base 2), v = (V base 1, v base 2). Let u = (1, 2), v = ( 2, - 1) and K = 3. Then the valued of <u, kv> is.....

(i) 4

(ii) 6

(iii) 18

(iv) None


19. Suppose that u, v "\isin" V are such that ||u|| = 2, ||u +v|| = 3 and ||u - v|| = 4. Then ||v|| is?

(i) 17/2

(ii) √17

(iii) Does not exist

(iv) None

20. For a given matrix A="\begin{bmatrix}
3 & 1 \\
 1 & 3
\end{bmatrix}", the matrix P that is orthogonally diagonalizes A is of the following matrices are diagonalisable

(i)P= "\begin{bmatrix}
1/√2 & 1/√2 \\
 1/√2 & - 1/√2
\end{bmatrix}"

(ii)P= "\begin{bmatrix}
 0 & 1 \\
 1 & 0 
\end{bmatrix}"

(iii)P="\begin{bmatrix}
-1/√2 & 1/√2 \\
 - 1/√2 & 1/√2
\end{bmatrix}"


1
Expert's answer
2022-01-11T17:06:00-0500

18.

"<u, v> =3u_ 1 v _ 1 + 2u _ 2 v _ 2"

"<u, kv>=3\cdot3\cdot2-2\cdot2\cdot3=18-12=6"

Answer: (ii) 6


19.

"||u \pm v|||^2=||u||^2\pm 2 <u, v>+||v||^2"

then:

"||u + v|||^2+||u - v|||^2=2(||u||^2+||v||^2)=9+16=25"

"||v||=\sqrt{25/2-4}=\sqrt{17/2}"

Answer: (iii) Does not exist


20.

"\begin{vmatrix}
 3-\lambda & 1 \\
 1 & 3-\lambda
\end{vmatrix}=0"


"\lambda^2-6\lambda+8=0"

"\lambda_1=2,\lambda_2=4"


"x+y=0\implies x=-y"

first eigenvector:

"\begin{pmatrix}
 x_1 \\
 -x_1
\end{pmatrix}"


"-x+y=0\implies x=y"

second eigenvector:

"\begin{pmatrix}
 x_2 \\
 x_2
\end{pmatrix}"


"P=\begin{pmatrix}
 x_2 &x_1\\
 x_2&-x_1
\end{pmatrix}"


Answer:

(i)P=Β "\begin{bmatrix}
1/√2 & 1/√2 \\
 1/√2 & - 1/√2
\end{bmatrix}"